# Optimizing Language Models for Human Preferences is a Causal Inference Problem

### v1

@@ 6.1 GPT-4 Annotation Validity
We use the online research platform Prolific [^5] https://www.prolific.co/ to conduct our human study. To avoid annotator fatigue, examples are annotated in batches of 20. We recruit a total of 30 annotators for an average of 3 annotators per example and a total of 600 annotations.
@@
Across three comparisons—human-human, majority vote-human, and human-GPT-4—we find that GPT-4 exhibits a similar or better level of agreement with human annotators as human annotators do with each other (Table 1). We conclude that GPT-4 is a reasonable surrogate for human annotators.
@@
Figure 1: CPO and DR-CPO win rates against $O$-RLHF, FT, and each other. A win rate exceeding 0.5 indicates that the named method outperforms the competing method with respect to the target outcome. Win rates are computed across 2000 pairs for each method combination.
@@
Figure 2: Impact of confounding (measured in win rate difference divided by 2) on $O$-RLHF, CPO, and DR-CPO. A negative impact indicates that confounding hurts the performance of the method. Win rates are computed across 600 pairs for each method combination.
6.2 Text Preferences
We report CPO and DR-CPO win rates against $O$-RLHF and FT (and against one another) in Figure 1. In Figure 2, we examine the impact of confounding on $O$-RLHF, CPO, and DR-CPO. Additional results are found in Appendix C.2.
@@ 6.2 Text Preferences
Outcome optimization (Figure 1). On the Hate Speech dataset, we observe that DR-CPO outperforms both the $O$-RLHF and FT baselines. Against both baselines, DR-CPO’s win rate is statistically significant at the 95% confidence level, with the lower bound of its 95% confidence intervals falling above 0.5. These results indicate that using DR-CPO, language models successfully learn human preferences for less hateful text from direct outcomes.
DR-CPO further appears to outperform CPO on the Hate Speech dataset, though its win rate is just shy of statistical significance. These results demonstrate the empirical benefits of double robustness, wherein a good $\widehat{P}^{R}$ provides bias-correction against a poorer $\widehat{g}$ (the former evidenced by DR-CPO’s relatively smaller win rate margin against CPO, and the latter evidenced by DR-CPO’s larger win rate margin against $O$-RLHF).
@@
On the Hong Kong dataset, we likewise observe that CPO outperforms both the $O$-RLHF and FT baselines. In this setting, CPO achieves statistically significant win rates against both baselines, with the lower bound of its 95% confidence intervals falling above 0.5.
Here, in contrast to the Hate Speech dataset, CPO also outperforms DR-CPO, which suggests that CPO can be very strong under conditions where $P^{R}$ is well controlled or can be estimated very well—as is the case for the Hong Kong dataset, where texts are not only randomly assigned to annotators but programmatically generated from random attributes. Furthermore, we note that outcome models trained on the Hong Kong dataset do not achieve good performance or generalization outside of the training data, possibly due to the artificial nature and relative homogeneity of the texts. Therefore, our results point to a conclusion that under conditions where the outcome model is particularly difficult to learn, CPO may enjoy empirical advantages over DR-CPO despite the theoretical robustness of the latter.
@@
Double robustness under confounding (Figure 2). Finally, on the Confounded dataset, we find that DR-CPO remains robust under confounding, while $O$-RLHF degrades significantly. (Vanilla CPO is not affected by confounding because it does not use an outcome model, but we train two separate models to account for randomness in the optimization process.) $O$-RLHF experiences a negative impact to win rate that is significantly lower than 0, while CPO experiences no impact and DR-CPO experiences a positive impact. Exploring this last result is an avenue for future work.
@@
These results further illustrate the doubly robust properties of DR-CPO and the shortcomings of outcome modeling approaches. Even under aggressive confounding, with a worst-case outcome model that has been trained on completely negated data, DR-CPO is not compromised, while $O$-RLHF is.
@@
We reiterate that because they are optimized on randomized experimental data, $\mathcal{D}_{R}$, CPO and DR-CPO are causal approaches. Taken together, our results constitute empirical evidence for a core theoretical strength—robustness to confounding—of an optimization framework that maintains the causal relationship between text and outcome.
@@ 7 Conclusion
In this paper, we explore language model optimization for human preferences from direct outcome datasets, in which each sample consists of a text and the reader’s numerical response. We first posit that language model optimization should be viewed as a causal problem to ensure that the model correctly learns the relationship between the text and the outcome, and we define conditions under which this causal relationship can be guaranteed. Following this, we introduce CPO, a method that solves an unbiased surrogate objective for the causal language optimization problem—and improve upon it with the doubly robust DR-CPO, which reduces the variance of the CPO objective while retaining provably strong guarantees on bias. Finally, we empirically demonstrate the effectiveness of (DR-)CPO in optimizing state-of-the-art LLMs for human preferences on direct outcome data, and we validate the robustness of DR-CPO under difficult confounding conditions. These theoretical contributions and results open the door to a wide range of data, human preferences, and optimization goals that language models can learn using CPO.
Several natural lines of future research follow from this work. For instance, (DR-)CPO may benefit empirically from exploration of entropy regularization techniques that are common in policy optimization. Additionally, future work may wish to extend DR-CPO to the paired completion data setting, as the bias guarantees and variance reduction of a doubly robust approach can also be useful for paired data.
@@ Acknowledgements.
This material is based upon work partially supported by Meta and the National Institutes of Health (awards R01MH125740, R01MH132225, and R21MH130767). Victoria Lin is supported by a Meta Research PhD Fellowship. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors, and no official endorsement should be inferred.
@@ A.2 Unbiasedness proofs and variance difference
[…]
@@
$(\ast)$ $=\mathbb{E}\left[\text{Var}\left(\frac{P^{f}(X)}{P^{R}(X)}(Y-\widehat{g}(X))\mid X\right)\right]+\text{Var}\left(\mathbb{E}\left[\frac{P^{f}(X)}{P^{R}(X)}(Y-\widehat{g}(X))\mid X\right]\right)$
@@ Appendix C Experiments
C.1 GPT-4 Win Rate Prompts
In this section, we include the inputs provided to GPT-4 for each dataset to obtain its preferences between texts generated by different methods. We use gpt-4-1106-preview from the OpenAI API. We adapt the prompt formats described in Rafailov et al. [2023]. The order of texts is random for each evaluation. For a (method, baseline) comparison, method corresponds to text A 50% of the time and text B 50% of the time.
C.1.1 Hate Speech prompt
@@
C.1.2 Hong Kong and Confounded prompt
@@ C.2 Additional Results
Table 2: CPO and DR-CPO win rates against $O$-RLHF, FT, and each other. A win rate exceeding 0.5 indicates that the named method outperforms the competing method with respect to the target outcome. Win rates are computed across 2000 pairs for each method combination.
@@
We include the full set of CPO and DR-CPO win rates against $O$-RLHF and FT and against each other (Table 2). We also include raw win rates from the confounding experiments, specifically win rates of unconfounded methods over confounded methods (Table 3). We briefly discuss comparisons that did not appear in the main body of the paper.
On the Hate Speech dataset, we observe that CPO—like DR-CPO—also outperforms both the $O$-RLHF and FT baselines. Against $O$-RLHF, CPO’s win rate is statistically significant at the 95% confidence level, while its win rate against FT falls slightly short of statistical significance.
On the Hong Kong dataset, we find that DR-CPO performs comparably to $O$-RLHF but falls short against the other methods. We attribute this to possible difficulty in learning the outcome model itself; this is further evidenced by the strong performance of CPO, which does not use an outcome model. As we mention in the main results, learning a strong outcome model on the Hong Kong dataset may be challenging, as its texts read somewhat artificially due to their programmatic construction.

### v2

@@ 6.1 GPT-4 Annotation Validity
We use the online research platform Prolific [^7] https://www.prolific.co/ to conduct our human study. To prevent annotator fatigue, examples are annotated in batches of 20. We recruit a total of 30 annotators for an average of 3 annotators per example and a total of 600 annotations.
@@
Across three comparisons—human-human, majority vote-human, and human-GPT-4—we find that GPT-4 exhibits a similar or better level of agreement with human annotators as human annotators do with each other (Table 1).
We conclude that GPT-4 is a reasonable surrogate for human annotators.
Dataset Method DR-CPO reward CPO reward OO-RLHF reward
Hate Speech FT 0.219 -0.024 0.242
Hate Speech OO-RLHF 0.244 -0.006 0.249
Hate Speech CPO 0.230 -0.012 0.242
Hate Speech DR-CPO 0.245 -0.005 0.250
Hong Kong FT 24.505 -0.256 24.761
Hong Kong OO-RLHF 22.256 -2.753 25.009
Hong Kong CPO 25.969 1.254 24.715
Hong Kong DR-CPO 25.162 0.269 24.893
Table 2: (Stabilized) expected reward under the DR-CPO, CPO, and OO-RLHF objectives. Expected reward under the DR-CPO objective is the most efficient estimate of the reward and therefore the most reliable metric. On the Hate Speech dataset, DR-CPO achieves the highest expected reward, while on the Hong Kong dataset, CPO achieves the highest reward.
@@
Figure 1: CPO and DR-CPO win rates against OO-RLHF, FT, and one another other. A win rate exceeding 0.5 indicates that the named method outperforms the competing method with respect to the target outcome. The error bars correspond to 95% confidence intervals, and asterisks (*) mean that the win rate difference between the two methods is statistically significant at the 95% confidence level.
@@
Figure 2: Impact of outcome model confounding (measured in win rate difference divided by 2) on OO-RLHF, CPO, and DR-CPO. A negative impact indicates that confounding hurts the performance of the method. The error bars correspond to 95% confidence intervals, and asterisks (*) mean that the win rate difference is statistically significant at the 95% confidence level.
6.2 Outcome Optimization
Over 2000 test prompts for each dataset, we evaluate CPO and DR-CPO win rates against baselines OO-RLHF and FT (and against one another). Under this evaluation framework, a win rate greater than 0.5 indicates that the method outperforms its competitor. We report our results in Figure 1. Additional results are found in Appendix C.3.
With the same 2000 prompts, we further evaluate expected reward [^8] Rewards are stabilized to prevent numerical underflow and overflow. under each optimization method. We reiterate that our optimization goal is to maximize the expected outcome or reward of texts generated by the language model. We report our results in Table 2.
@@ 6.2 Outcome Optimization
On the Hate Speech dataset, we observe that DR-CPO outperforms both the OO-RLHF and FT baselines with respect to win rate. Against both baselines, DR-CPO’s win rate is statistically significant at the 95% confidence level, with the lower bound of its 95% confidence intervals falling above 0.5. DR-CPO also achieves the highest expected reward across all objectives. These results indicate that using DR-CPO, language models successfully learn human preferences for less hateful text from direct outcomes—and do so better than competing baselines.
DR-CPO additionally appears to outperform CPO on the Hate Speech dataset, though its win rate is just shy of statistical significance. This result supports our theoretical result (Proposition 4.5) that the addition of an outcome model to the CPO objective can improve optimization by reducing variance. Moreover, the success of DR-CPO over OO-RLHF provides further evidence for the benefits of the doubly robust formulation, wherein a good $P^{R}$ can help to correct bias from an outcome modeling approach.
@@
On the Hong Kong dataset, we likewise observe that CPO outperforms both the OO-RLHF and FT baselines with respect to both win rate and expected reward under the DR-CPO objective. In this setting, CPO achieves statistically significant win rates against both baselines, with the lower bound of its 95% confidence intervals falling above 0.5.
Here, in contrast to the Hate Speech dataset, CPO also outperforms DR-CPO, which suggests that CPO can be very strong under conditions where $P^{R}$ is well controlled or can be estimated very well—as is the case for the Hong Kong dataset, where texts are not only randomly assigned to annotators but programmatically generated from random attributes. An additional factor may be the difficulty of modeling the complex continuous outcome of the Hong Kong dataset from its artificially constructed, homogeneous texts. Under such conditions, an outcome model may introduce noise rather than help with bias correction (see Proposition 4.5 for a discussion of the objectives’ precision), and so CPO may enjoy empirical advantages over DR-CPO. The challenge of outcome modeling is additionally reflected in the underperformance of OO-RLHF relative to even FT, providing evidence for how outcome modeling alone can be insufficient for language model optimization.
6.3 Double Robustness Under Confounding
In Figure 2, we examine the impact of outcome model confounding on OO-RLHF, CPO, and DR-CPO. For our confounded condition, we train both the DR-CPO and OO-RLHF outcome models on the Confounded dataset. As DR-CPO also uses randomized data in its IPW term, we optimize this portion of the objective on the (unconfounded) Hong Kong dataset. For each method, we then compute win rate between a model that has been optimized under these confounding conditions and a model that has been optimized on unconfounded data using the same method.
CPO is constructed to use only randomized data and does not have an outcome model, so it is exclusively optimized on the (unconfounded) Hong Kong dataset. However, we compare two separately optimized models to account for randomness in the optimization process.
Win rates are computed across 600 pairs for each method. Additional results are found in Appendix C.3.
@@ 6.3 Double Robustness Under Confounding
In these experiments, we find that DR-CPO remains robust under confounding in the outcome model, while OO-RLHF degrades significantly. OO-RLHF experiences an impact to win rate that is significantly negative, while CPO experiences no impact (as expected, because it does not use an outcome model) and DR-CPO experiences a positive impact. Exploring this last result is an avenue for future work.
@@
These results highlight one of the core strengths of DR-CPO and one of the core shortcomings of outcome modeling approaches. As long as a small crowdsourced dataset is available to act as $P^{R}$, DR-CPO can use large amounts of confounded data to learn its outcome model and still retain its unbiasedness guarantees—all while reducing variance relative to CPO. An exclusively outcome model-based approach like OO-RLHF, on the other hand, becomes biased under these conditions. Even under aggressive confounding, with a worst-case outcome model that has been trained on completely negated data, DR-CPO is not compromised, while OO-RLHF is.
@@
We reiterate that because they are optimized (at least partially) on randomized experimental data, CPO and DR-CPO are causal approaches. Taken together, our results emphasize the importance of an optimization framework that maintains the causal relationship between text and outcome.
@@ 7 Conclusion
In this paper, we explore language model optimization for human preferences from direct outcome datasets, in which each sample consists of a text and the reader’s numerical response. We first posit that language model optimization should be viewed as a causal problem to ensure that the model correctly learns the relationship between the text and the outcome, and we define conditions under which this causal relationship can be guaranteed. Following this, we introduce CPO, a method that solves an unbiased surrogate objective for the causal language optimization problem—and improve upon it with the doubly robust DR-CPO, which reduces the variance of the CPO objective while retaining provably strong guarantees on bias. Finally, we empirically demonstrate the effectiveness of (DR-)CPO in optimizing state-of-the-art LLMs for human preferences on direct outcome data, and we validate the robustness of DR-CPO under difficult confounding conditions.
To our knowledge, our work is the first to approach LLM optimization as a causal inference problem—from which an importance-weighted solution naturally follows—as well as the first to propose a doubly robust methodology for LLM optimization. These theoretical contributions and results open the door to a wide range of data, human preferences, and optimization goals that language models can learn using (DR-)CPO.
Several natural lines of future research follow from this work. For instance, (DR-)CPO may benefit empirically from entropy regularization techniques that are common in policy optimization. Additionally, the performance gains in our evaluations may have been limited by the small proportion of parameters updated under LoRA, and so it would be useful to explore whether experiments with commercial-grade computational resources have the potential to yield much larger improvements. Finally, future work may wish to extend DR-CPO to the paired completion data setting, as the bias guarantees and variance reduction of a doubly robust approach can also be useful for paired data.
@@ Acknowledgements.
This material is based upon work partially supported by the National Institutes of Health (awards R01MH125740, R01MH132225, and R21MH130767). Victoria Lin is supported by a Meta Research PhD Fellowship. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors, and no official endorsement should be inferred.
@@ A.2 Unbiasedness proofs and variance difference
$=\frac{1}{n}\underbrace{\text{Var}\left(\frac{P^{f}(X)}{P^{R}(X)}(Y-\widehat{g}(X))\right)}_{(\ast)}+\frac{1}{m}\text{Var}\left(\frac{P^{f}(\widetilde{X})}{P^{f^{0}}(\widetilde{X})}\widehat{g}(\widetilde{X})\right)$
@@
(($\ast$)( ∗ ) $=\mathbb{E}\left[\text{Var}\left(\frac{P^{f}(X)}{P^{R}(X)}(Y-\widehat{g}(X))\mid X\right)\right]+\text{Var}\left(\mathbb{E}\left[\frac{P^{f}(X)}{P^{R}(X)}(Y-\widehat{g}(X))\mid X\right]\right)$
@@ Appendix C Experiments
C.1 Model details and hyperparameters
Our language models were implemented using the HuggingFace transformers library (version 4.32.1), with Llama 2 7B weights stored locally. For each language model optimization method, we trained for 3 epochs with a batch size of 4, maximum gradient norm of 1.0, and learning rate of $5\times 10^{-5}$, with all other hyperparameters set to Trainer class defaults. We used the HuggingFace peft library for LoRA and set our LoraConfig parameters as follows: rank 8, alpha 8, dropout probability 0.1, with all other parameters remaining as default. Custom loss functions for CPO, DR-CPO, and OO-RLHF were implemented using PyTorch (torch version 2.0.1).
For the Hong Kong dataset outcome model, we used scikit-learn (version 1.3.0) to implement our linear regression. All hyperparameters were set to their default values. Empath features were extracted using the empath library. [^9] https://github.com/Ejhfast/empath-client
C.2 GPT-4 Win Rate Prompts
Here, we include the inputs provided to GPT-4 for each dataset to obtain its preferences between texts generated by different methods. We accessed GPT-4 through the OpenAI API (gpt-4-1106-preview). Prompt formats are adapted from those described in Rafailov et al. [2023]. The order of texts is random for each evaluation. For a (method, baseline) comparison, method corresponds to text A 50% of the time and text B 50% of the time.
C.2.1 Hate Speech prompt
@@
C.2.2 Hong Kong and Confounded prompt
@@ C.3 Additional Results
Table 3: CPO and DR-CPO win rates against OO-RLHF, FT, and each other. A win rate exceeding 0.5 indicates that the named method outperforms the competing method with respect to the target outcome. Win rates are computed across 2000 pairs for each method combination.
@@
We include the full set of CPO and DR-CPO win rates against OO-RLHF and FT and against each other (Table 3). We also include raw win rates from the confounding experiments, specifically win rates of unconfounded methods over confounded methods (Table 4). We briefly discuss comparisons that did not appear in the main body of the paper.
On the Hate Speech dataset, we observe that CPO—like DR-CPO—also outperforms both the OO-RLHF and FT baselines. Against OO-RLHF, CPO’s win rate is statistically significant at the 95% confidence level, while its win rate against FT falls slightly short of statistical significance.
On the Hong Kong dataset, we find that DR-CPO performs comparably to OO-RLHF but falls short against the other methods. We attribute this to possible difficulty in learning the outcome model itself; this is further evidenced by the strong performance of CPO, which does not use an outcome model. As we mention in the main results, learning a strong outcome model on the Hong Kong dataset may be challenging, as its texts read somewhat artificially due to their programmatic construction.